# On the Power of the Weisfeiler-Leman Test for Graph Motif Parameters

### v1

@@ B.1 Proof of Theorem 15
Let $H$ be the disjoint union of $\mathsf{CCFI}(G_{1})$ and $J$, and let $H^{\prime}$ be the disjoint union of $\mathsf{CCFI}^{\times}(G_{1})$ and $J$. We will show that $H\equiv_{\mathcal{F}}H^{\prime}$ and $\mathsf{homs}(G,H)\neq\mathsf{homs}(G,H^{\prime})$. For the former, suppose $F\in\mathcal{F}$ and connected. By Condition (1), there is no weak oddomorphism from $F$ to $G$, and thus $\mathsf{homs}(F,\mathsf{CCFI}(G_{1}))=\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G_{1}))$ according to Theorem 22. Because $F$ is connected each homomorphism into $H$ is either a homomorphism into $J$ or into $\mathsf{CCFI}(G_{1})$, i.e., $\mathsf{homs}(F,H)=\mathsf{homs}(F,\mathsf{CCFI}(G_{1}))+homs(F,J)$. The same holds for $H^{\prime}$, i.e., $\mathsf{homs}(F,H^{\prime})=\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G_{1}))+\mathsf{homs}(F,J)$. Since both terms in both sums are equal we see that $\mathsf{homs}(F,H)=\mathsf{homs}(F,H^{\prime})$. Since this holds for all connected $F\in\mathcal{F}$, by Condition (2), it holds for all $F\in\mathcal{F}$.
@@
B.2 On Vertex Labels
Our discussion so far has focused only on edge-labeled graphs, i.e., labeled graphs where the vertex labeling function $\lambda$ maps all vertices to the same label. In this section, we show that the general case with arbitrary vertex labels can be reduced to the edge labeled case.
Let $G=(V,E,\lambda,\kappa)$ be a labeled graph. Define $\mathsf{EL}(G)$ as the labeled graph $(V,E,\lambda^{\prime},\kappa^{\prime})$ where $\lambda^{\prime}(v)=1$ for all $v\in V$ and $\kappa^{\prime}(\{a,b\})=(\lambda(a),\lambda(b),\kappa(\{a,b\}),b\}\in E$ and $\lambda(a)\leq_{lex}\lambda(b)$.
Proposition 25.
Let $F,G$ be labeled graphs. Then $\phi\in\mathsf{Hom}(F,G)$ if and only if $\phi\in\mathsf{Hom}(\mathsf{EL}(F),\mathsf{EL}(G))$
Proof.
The implication left to right is straightforward. For the other direction, suppose $\phi\in\mathsf{Hom}(\mathsf{EL}(F),\mathsf{EL}(G))$. It is clear that for any edge $e\in E(F)$, $\phi(e)\in E(G)$ as the sets are not changed by the $\mathsf{EL}$ function. Furthermore, since $\kappa^{\prime}(e)=\kappa^{\prime}(\phi(e))$, the tuples must agree on the third component, i.e., $\kappa(e)=\kappa(\phi(e))$. Now suppose $e=\{u,v\}$ and, w.l.o.g., $\lambda(u)\leq\lambda(v)$. Then the first component of $\kappa^{\prime}(e)$ is $\lambda(u)$, and therefore so is the first component of $\kappa^{\prime}(\phi(e))$, i.e., $\lambda(u)=\lambda(\phi(u))$. Similarly, from the second component we get $\lambda(v)=\lambda(\phi(v))$. ∎
If we encode a labeled graph as an edge-labeled graph, it is clear that this encoding can be reverted again. We will need to observe that this “invertibility” carries over also to the $\mathsf{CCFI}$ construction.
Proposition 26.
Let $G$ be an labeled graph. For every $U\subseteq V(G)$, there exists a labeled graph $H$ such that $\mathsf{EL}(H)=\mathsf{CCFI}(\mathsf{EL}(G),U)$.
Proof.
Let $H$ be the labeled graph with the vertices and edges of $\mathsf{CCFI}(\mathsf{EL}(G,U))$. For the labels of $H$, set $\lambda_{H}((v,S))=\lambda_{G}(v)$ and $\kappa_{H}(\{(v,S),(u,T)\})=\kappa_{G}(\{v,u\})$, where the subscripts distinguish the labeling functions of $G$ and $H$. Any edge between $(v,S)$ and $(u,T)$ with $\lambda_{G}(v)\leq_{lex}\lambda_{G}(u)$ in $\mathsf{EL}(H)$ will thus be labeled as $(\lambda_{G}(v),\kappa_{G}(\{v,u\},\lambda_{G}(u))$. This is precisely the label of the edge $\{v,u\}$ in $\mathsf{EL}(G)$, and thus inherited for $\{(v,S),(u,T)\}$ in the construction of $\mathsf{CCFI}(G,U)$. That is, $\mathsf{EL}(H)$ and $\mathsf{CCFI}(\mathsf{EL}(G),U)$ have the same vertices, edges, and labels. ∎
We will also want a similar invertibility from the graph chosen from Lemma 24.
Lemma 27.
Fix integer $n\geq 1$ and vertex- and edge-label alphabets $\Sigma$ and $\Delta$. There is a connected labeled graph $G$ with vertex- and edge-labels $\Sigma$ and $\Delta$ such that for every labeled graph $F$ with $n$ vertices and vertex- and edge-labels $\Sigma$ and $\Delta$, $\mathsf{homs}(\mathsf{EL}(F),\mathsf{EL}(G))>0$.
Proof.
Let $\mathcal{C}$ be the set of all (up to isomorphism) labeled $n$ vertex cliques over vertex- and edge-label alphabets $\Sigma$ and $\Delta$. The desired graph $G$ is then constructed by taking the disjoint union of $\mathcal{C}$, adding a new vertex $v$ that is adjacent to one vertex of every $C\in\mathcal{C}$. The labels of $v$ and its incident edges can be chosen arbitrarily. Clearly, there is some $C\in\mathcal{C}$ such that for any considered $F$, $\mathsf{homs}(F,C)>0$ and thus also $\mathsf{homs}(\mathsf{EL}(F),\mathsf{EL}(C))>0$. Since $C$ is a subgraph of $G$, $\mathsf{EL}(C)$ is a subgraph of $\mathsf{EL}(G)$. ∎
Proof of Lemma 3.
Let $G$ be a labeled graph of treewidth greater than $k$. By inspection of the proof of Theorem 15 and Lemma 17, there exist edge-labeled graphs $H,H^{\prime}$ such that $H\equiv_{\mathcal{ET}_{k}}H^{\prime}$, $\mathsf{homs}(\mathsf{EL}(G),H)\neq\mathsf{homs}(\mathsf{EL}(G),H^{\prime})$, $H$ is the disjoint union of $\mathsf{CCFI}(\mathsf{EL}(G_{1}))$ and some $J$, and $H^{\prime}$ is the disjoint union of $\mathsf{CCFI}^{\times}(\mathsf{EL}(G_{1}))$ and $J$, where $G_{1}$ is a maximal connected component of $G$ with maximal treewidth.
By Proposition 26, there are labeled graphs $X,X^{\prime}$ such that $\mathsf{EL}(X)=\mathsf{CCFI}(\mathsf{EL}(G_{1})$ and $\mathsf{EL}(X^{\prime})=\mathsf{CCFI}^{\times}(\mathsf{EL}(G_{1}))$. By Lemma 27, we can assume w.l.o.g, that there is a labeled graph $Y$ such that $\mathsf{EL}(Y)=J$. Let $Z$ be the disjoint union of $X$ and $Y$, and $Z^{\prime}$ be the disjoint union of $X^{\prime}$ and $Y$. We see that $\mathsf{EL}(Z)=H$ and $\mathsf{EL}(Z^{\prime})=H^{\prime}$ and thus
$\mathsf{homs}(G,Z)=\mathsf{homs}(\mathsf{EL}(G),H)\neq\mathsf{homs}(\mathsf{EL}(G),H^{\prime})=\mathsf{homs}(G,Z^{\prime})$
where the equalities are by Proposition 25.
We also know that for every edge-labeled graph $F^{*}\in\mathcal{ET}_{k}$, $\mathsf{homs}(F^{*},H)=\mathsf{homs}(F^{*},H^{\prime})$. Thus also for every labeled graph $F\in\mathcal{LT}_{k}$, $\mathsf{homs}(F,Z)=\mathsf{homs}(\mathsf{EL}(F),H)=\mathsf{homs}(\mathsf{EL}(F),H^{\prime})=\mathsf{homs}(F,Z^{\prime})$. In other words $Z\equiv_{\mathcal{LT}_{k}}Z^{\prime}$ and the proof is complete. ∎
B.3 Lemma 5
@@
As a first step, we consider the logic $\mathcal{C}^{L}_{k+1}$ of first-order formulas with $k+1$ variables and counting quantifiers [^4] For an extensive definition of first-order logic with counting quantifiers refer to Cai et al. (1992). over the signature that contains unary relation symbols $U_{\sigma}$ for each label $\sigma$ in the vertex label alphabet $\Sigma$, and binary relation symbols $E_{\delta}$ for each label $\delta$ in the edge label alphabet $\Delta$. We add the superscript $L$ here to distinguish from the usual use of $\mathcal{C}_{k+1}$ for unlabeled graphs. Cai et al. (1992) famously showed that for graphs $G,H$ we have $G\models\varphi\iff H\models\varphi$ for all $\varphi\in\mathcal{C}_{k+1}$ if and only if $G\equiv_{kWL}H$. It is folklore that this also holds in the setting of labeled graphs. We in fact only need one direction here, namely if $G\not\equiv_{kWL}H$, then there is a $\varphi\in\mathcal{C}^{L}_{k+1}$ such that $G\models\varphi$ and $H\not\models\varphi$. It is straightforward to verify that the original argument for this direction is unaffected by labels (Cai et al., 1992, Theorem 5.4, case $\neg 1\Rightarrow\neg 2$).
@@
We first recall the key definitions from (Dvorák, 2010). A $k$-marked [^5] What we call $k$-marked here is referred to as $k$-labeled in the original which could clearly clash with our use of labeled graphs. labeled graph $G$ is a graph with a partial function $\mathsf{mark}_{G}:[k]\to V(G)$. The set of active markings $M_{G}$ is the subset of $[k]$ for which $\mathsf{mark}_{G}$ is defined. Suppose that $G,H$ are $k$-marked labeled graphs with $M_{H}\subseteq M_{H}$, we define $\mathsf{homs}(G,H)$ as the number of those homomorphisms $\phi$ that also preserve markings, i.e., where $\phi(\mathsf{mark}_{G}(i))=\mathsf{mark}_{H}(i)$ for each active marking $i\in M_{G}$. A $k$-marked labeled quantum graph $G$ is a finite linear combination with real coefficients of $k$-marked labeled quantum graphs. We require all graphs in the linear combination to have the same set of active markings, which we refer to as $M_{G}$. The function $\mathsf{homs}$ is extended to the case where the first argument is a $k$-marked labeled quantum graph $G=\sum_{i}^{\ell}\alpha_{i}G_{i}$ as $\mathsf{homs}(G,H)=\sum_{i}^{\ell}\alpha_{i}\mathsf{homs}(G_{i},H)$. The treewidth of a quantum graph $\sum_{i}\alpha_{i}G_{i}$ is $\max\{tw(G_{i})\mid\alpha_{i}\neq 0\}$ [^6] Homomorphisms from quantum graphs can be seen as an alternative representation of graph motif parameters. Because of the two distinct natures of their uses in this paper and the disjoint connections to prior work we have decided to not unify the presentation of these two concepts here..
@@
As discussed above, we have that for labeled graphs $G,H$, $G\not\equiv_{kWL}H$ implies the existence of a $\phi\in\mathcal{C}^{L}_{k+1}$ such that $G\models\varphi$ and $H\not\models\varphi$. By Lemma 28, there is a $k$-marked labeled quantum graph $F$ with treewidth at most $k$ that models $\varphi$. Hence $\sum_{i}\alpha_{i}\mathsf{homs}(F_{i},G)=1\neq 0=\sum_{i}\alpha_{i}\mathsf{homs}(F_{i},H)$, which means there must be some $i$ such that $\mathsf{homs}(F_{i},H)\neq\mathsf{homs}(F_{i},H)$ (recall that the linear combination is always finite). Since $F$ has treewidth at most $k$, $F_{i}\in\mathcal{LT}_{k}$ and therefore $G\not\equiv_{\mathcal{LT}_{k}}H$. ∎
B.4 On Lemma 4 for Labeled Graphs
@@ Appendix C Proof of Theorem 8
This section will make some slight departures from the terminology used in the main body of the paper. For a more focused presentation, we will first discuss the case for unlabeled graphs and afterwards show how to additionally handle labels afterwards. In the main body, the $\mathsf{spasm}$ of a graph was defined as the set of all homomorphic images, but here it will be simpler to take an alternative (equivalent) perspective. A quotient $G/\tau$ of a graph $G$ is obtained by taking a partition $\tau$ of $V(G)$ and constructing the graph like $G$ but with all vertices in the same block of $\tau$ identified. That is, the vertices of $G$ are the blocks of $\tau$, and the incidence of a vertex in $G\/sigma$ is the incidence of all vertices in the corresponding block of the partition. It is a standard observation that the set of all quotients of $G$ without self-loops is precisely $\mathsf{spasm}(G)$ (see, e.g., Curticapean et al. (2017)). It will be convenient to also consider the set of all quotients of $G$, i.e., including those with self-loops, which we will refer to as $\mathsf{spasm}^{\circ}$.
@@
Checking whether graph $H$ is a minor of graph $G$ via an MSO formula is a standard construction. Interestingly, to check whether $H$ is a minor of some quotient of $G$ is simpler, at least in terms of formulating the question in MSO. The key insight here is that the second condition of minor models can effectively be ignored, since there will always be a quotient where all vertices of $G[f(v)]$ are identified. Instead it is enough to guarantee that the image is non-empty. In concrete terms, to check whether $H$ is a minor of a graph in $\mathsf{spasm}^{\circ}(G)$ we will use the following MSO formula.
@@
We can resolve the possibility of self-loops with a simple observations about minors of quotient graphs. A self-loop occurs in quotient $G/\tau$ if there is a block $B$ of $\tau$ that contains adjacent vertices $v,u$. W.l.o.g., assume for now that $v,u$ are the only adjacent vertices in $B$. If $H$ is a minor of $G/\tau$, it is also a minor of $G/\tau^{\prime}$ where $\tau^{\prime}$ is like $\tau$ but with $u$ as a singleton rather than in $B$. The vertex for $\{u\}$ in $G/\tau^{\prime}$ will be adjacent to the vertex for block $B^{\prime}=B\setminus\{u\}$. Contracting the vertices for $B^{\prime}$ and $\{u\}$ will produce exactly $G/\tau$ without the self-loop at $B$. Iterating this idea shows that $G/\tau$ without self-loops is a minor of some loop-free quotient of $G$.
@@
We first check whether $tw(G)\leq k-1$. It is well known that this can be decided in linear time for fixed $k$ Bodlaender (1996). If not, then $tw(G)\geq k$ and the algorithm returns true.
The property of having treewidth at most $k-1$ is closed under graph minors. By Theorem 31 there exists a finite set $\mathcal{F}$ of forbidden minors for class of graphs with treewidth at most $k-1$. Furthermore, the set $\mathcal{F}$ depends only on the parameter $k$. Thus, we have that a graph $H$ has treewidth $\geq k$ iff it does not have treewidth $\leq k-1$ iff it has some some graph of $\mathcal{F}$ as a minor. Combining with Proposition 30, we can therefore decide whether the treewidth in the spasm is at least $k$, by checking for every $F\in\mathcal{F}$ whether $F$ is a minor of any quotient of $G$. According to Lemma 29, this is equivalent to checking whether the formula
@@
is satisfied by $G$. By Courcelle’s Theorem (Courcelle, 1990) we can decide $G\models\phi$ in time $f(k)\cdot|G|$ where $k$ is the treewidth of $G$. ∎
As a final note we wish to note that this result is primarily of theoretical interest. For most practical pattern sizes naive enumeration of the $\mathbf{spasm}$ and checking the treewidth individually with state of the art systems for treewidth computation is feasible. The bottleneck in practice is the size of the pattern, as enumerating the $\mathsf{spasm}$ via naive methods, e.g., enumerating all partitions, (see Sequence A000110 in the On-Line Encyclopedia of Integer Sequences (OEIS Foundation Inc., 2023)).
@@
We finally discuss the necessary changes for labeled graphs in the above argument, from which Theorem 8 then follows immediately. By definition, treewidth of labeled graphs is unaffected by the labels. As such we can can simply ignore labels in checking for minors. Where the labels make a difference however, is in the set $\mathsf{spasm}$ itself. Not all loop-free quotients are homomorphic images in the labeled case, but rather we require an additional restriction.

### v3

@@ B.1 Proof of Theorem 15
Let $H$ be the disjoint union of $\mathsf{CCFI}(G_{1})$ and $J$, and let $H^{\prime}$ be the disjoint union of $\mathsf{CCFI}^{\times}(G_{1})$ and $J$. We will show that $H\equiv_{\mathcal{F}}H^{\prime}$ and $\mathsf{homs}(G,H)\neq\mathsf{homs}(G,H^{\prime})$. For the former, suppose $F\in\mathcal{F}$ and connected. By Condition (1), there is no weak oddomorphism from $F$ to $G$, and thus $\mathsf{homs}(F,\mathsf{CCFI}(G_{1}))=\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G_{1}))$ according to Theorem 22. Because $F$ is connected each homomorphism into $H$ is either a homomorphism into $J$ or into $\mathsf{CCFI}(G_{1})$, i.e., $\mathsf{homs}(F,H)=\mathsf{homs}(F,\mathsf{CCFI}(G_{1}))+\mathsf{homs}(F,J)$. The same holds for $H^{\prime}$, i.e., $\mathsf{homs}(F,H^{\prime})=\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G_{1}))+\mathsf{homs}(F,J)$. Since both terms in both sums are equal we see that $\mathsf{homs}(F,H)=\mathsf{homs}(F,H^{\prime})$. Since this holds for all connected $F\in\mathcal{F}$, by Condition (2), it holds for all $F\in\mathcal{F}$.
@@
B.2 Lemma 5
@@
As a first step, we consider the logic $\mathcal{C}^{L}_{k+1}$ of first-order formulas with $k+1$ variables and counting quantifiers [^4] For a detailed definition of first-order logic with counting quantifiers refer to Cai et al. (1992). over the signature that contains unary relation symbols $U_{\sigma}$ for each label $\sigma$ in the vertex label alphabet $\Sigma$, and binary relation symbols $E_{\delta}$ for each label $\delta$ in the edge label alphabet $\Delta$. We add the superscript $L$ here to distinguish from the usual use of $\mathcal{C}_{k+1}$ for unlabeled graphs. Cai et al. (1992) famously showed that for graphs $G,H$ we have $G\models\varphi\iff H\models\varphi$ for all $\varphi\in\mathcal{C}_{k+1}$ if and only if $G\equiv_{kWL}H$. It is folklore that this also holds in the setting of labeled graphs. We in fact only need one direction here, namely if $G\not\equiv_{kWL}H$, then there is a $\varphi\in\mathcal{C}^{L}_{k+1}$ such that $G\models\varphi$ and $H\not\models\varphi$. It is straightforward to verify that the original argument for this direction is unaffected by labels (Cai et al., 1992, Theorem 5.4, case $\neg 1\Rightarrow\neg 2$).
@@
We first recall the key definitions from (Dvorák, 2010). A $k$-marked [^5] What we call $k$-marked here is referred to as $k$-labeled in (Dvorák, 2010) and much of the literature. Unfortunately, this would clash with the other with our (also common) terminology for graphs with vertex and edge labels. labeled graph $G$ is a graph with a partial function $\mathsf{mark}_{G}:[k]\to V(G)$. The set of active markings $M_{G}$ is the subset of $[k]$ for which $\mathsf{mark}_{G}$ is defined. Suppose that $G,H$ are $k$-marked labeled graphs with $M_{H}\subseteq M_{H}$, we define $\mathsf{homs}(G,H)$ as the number of those homomorphisms $\phi$ that also preserve markings, i.e., where $\phi(\mathsf{mark}_{G}(i))=\mathsf{mark}_{H}(i)$ for each active marking $i\in M_{G}$. A $k$-marked labeled quantum graph $G$ is a finite linear combination with real coefficients of $k$-marked labeled quantum graphs. We require all graphs in the linear combination to have the same set of active markings, which we refer to as $M_{G}$. The function $\mathsf{homs}$ is extended to the case where the first argument is a $k$-marked labeled quantum graph $G=\sum_{i}^{\ell}\alpha_{i}G_{i}$ as $\mathsf{homs}(G,H)=\sum_{i}^{\ell}\alpha_{i}\mathsf{homs}(G_{i},H)$. The treewidth of a quantum graph $\sum_{i}\alpha_{i}G_{i}$ is $\max\{tw(G_{i})\mid\alpha_{i}\neq 0\}$ [^6] Homomorphisms from quantum graphs can be seen as an alternative representation of graph motif parameters. Because of the two distinct natures of their uses in this paper and the disjoint connections to prior work we have decided to not unify the presentation of these two concepts here..
@@
As discussed above, we have that for labeled graphs $G,H$, $G\not\equiv_{kWL}H$ implies the existence of a $\phi\in\mathcal{C}^{L}_{k+1}$ such that $G\models\varphi$ and $H\not\models\varphi$. By Lemma 25, there is a $k$-marked labeled quantum graph $F$ with treewidth at most $k$ that models $\varphi$. Hence, $\sum_{i}\alpha_{i}\mathsf{homs}(F_{i},G)=1\neq 0=\sum_{i}\alpha_{i}\mathsf{homs}(F_{i},H)$, which means there must be some $i$ such that $\mathsf{homs}(F_{i},H)\neq\mathsf{homs}(F_{i},H)$ (recall that the linear combination is always finite). Since $F$ has treewidth at most $k$, $F_{i}\in\mathcal{LT}_{k}$ and therefore $G\not\equiv_{\mathcal{LT}_{k}}H$. ∎
B.3 On Lemma 4 for Labeled Graphs
@@ Appendix C Proof of Theorem 8
This section will make some slight departures from the terminology used in the main body of the paper. For a more focused presentation, we will first discuss the case for unlabeled graphs and afterwards show how to additionally handle labels afterwards. In the main body, the $\mathsf{Spasm}$ of a graph was defined as the set of all homomorphic images, but here it will be simpler to take an alternative (equivalent) perspective. A quotient $G/\tau$ of a graph $G$ is obtained by taking a partition $\tau$ of $V(G)$ and constructing the graph like $G$ but with all vertices in the same block of $\tau$ identified. That is, the vertices of $G$ are the blocks of $\tau$, and the incidence of a vertex in $G/\sigma$ is the incidence of all vertices in the corresponding block of the partition. It is a standard observation that the set of all quotients of $G$ without self-loops is precisely $\mathsf{Spasm}(G)$ (see, e.g., Curticapean et al. (2017)). It will be convenient to also consider the set of all quotients of $G$, i.e., including those with self-loops, which we will refer to as $\mathsf{Spasm}^{\circ}$.
@@
Checking whether graph $H$ is a minor of graph $G$ via an MSO formula is a standard construction. Interestingly, to check whether $H$ is a minor of some quotient of $G$ is simpler, at least in terms of formulating the question in MSO. The key insight here is that the second condition of minor models can effectively be ignored, since there will always be a quotient where all vertices of $G[f(v)]$ are identified. Instead, it is enough to guarantee that the image is non-empty. In concrete terms, to check whether $H$ is a minor of a graph in $\mathsf{Spasm}^{\circ}(G)$ we will use the following MSO formula.
@@
We can resolve the possibility of self-loops with a simple observation about minors of quotient graphs. A self-loop occurs in quotient $G/\tau$ if there is a block $B$ of $\tau$ that contains adjacent vertices $v,u$. W.l.o.g., assume for now that $v,u$ are the only adjacent vertices in $B$. If $H$ is a minor of $G/\tau$, it is also a minor of $G/\tau^{\prime}$ where $\tau^{\prime}$ is like $\tau$ but with $u$ as a singleton rather than in $B$. The vertex for $\{u\}$ in $G/\tau^{\prime}$ will be adjacent to the vertex for block $B^{\prime}=B\setminus\{u\}$. Contracting the vertices for $B^{\prime}$ and $\{u\}$ will produce exactly $G/\tau$ without the self-loop at $B$. Iterating this idea shows that $G/\tau$ without self-loops is a minor of some loop-free quotient of $G$.
@@
We first check whether $tw(G)\leq k-1$. It is well known that this can be decided in linear time for fixed $k$ (Bodlaender, 1996). If not, then $tw(G)\geq k$ and the algorithm returns true.
The property of having treewidth at most $k-1$ is closed under graph minors. By Theorem 28 there exists a finite set $\mathcal{F}$ of forbidden minors for class of graphs with treewidth at most $k-1$. Furthermore, the set $\mathcal{F}$ depends only on the parameter $k$. Thus, we have that a graph $H$ has treewidth $\geq k$ iff it does not have treewidth $\leq k-1$ iff it has some graph of $\mathcal{F}$ as a minor. Combining with Proposition 27, we can therefore decide whether the treewidth in the spasm is at least $k$, by checking for every $F\in\mathcal{F}$ whether $F$ is a minor of any quotient of $G$. According to Lemma 26, this is equivalent to checking whether the formula
@@
is satisfied by $G$. By Courcelle’s Theorem (Courcelle, 1990) we can decide $G\models\phi$ in time $f(k,\varphi)\cdot|G|$ where $k$ is the treewidth of $G$. Since $\varphi$ depends only on $k$ this is equivalent to $f(k)\cdot|G|$ and hence polynomial for fixed $k$. ∎
As a final note we wish to point out that this result is primarily of theoretical interest. For most practical pattern sizes, naive enumeration of $\mathsf{Spasm}$ and checking the treewidth individually with state-of-the-art systems for treewidth computation is feasible. The bottleneck in practice is the size of the pattern, as enumerating the contents of $\mathsf{Spasm}$ via naive methods, e.g., enumerating all partitions, becomes challenging above 11 vertices (see Sequence A000110 in the On-Line Encyclopedia of Integer Sequences (OEIS Foundation Inc., 2023)).
@@
We finally discuss the necessary changes for labeled graphs in the above argument, from which Theorem 8 then follows immediately. By definition, treewidth of labeled graphs is unaffected by the labels. As such we can simply ignore labels in checking for minors. Where the labels make a difference however, is in the set $\mathsf{Spasm}$ itself. Not all loop-free quotients are homomorphic images in the labeled case, but rather we require an additional restriction.